OpenStudy (anonymous):

A science teacher has a supply of 50% sugar solution and a supply of 80% sugar solution. How much of each solution should the teacher mix together to get 150 mL of 60% sugar solution? a. 70 mL of the 50% solution and 35mL of the 80% solution. n. 35mL of the 50% solution and 70mL of the 80% solution. c. 70mL of the 50% solution and 70mL of the 80% solution. d. 35mL of the 50% solution and 35mL of the 80% solution.

OpenStudy (studygurl14):

mLs of 50% sugar solution = x mLs of 80% solution = y x + y = 150 mL 50% + 80% = 60% Put them together and change the %'s into decimals... x(0.5) + y(0.8) = 150(0.6) Now rearrange this equation: x + y = 150 x = 150 - y y = 10 -x What now?

OpenStudy (anonymous):

umm......